OpenStudy (anonymous):

Hi I need help evaluating an integral, it is: 0 to 1 squareroot(1+e^-x)/e^x dx, I'm not sure if I substitute twice or just once?

terenzreignz (terenzreignz):

I find staring at it for a while helps... \[\Large \int\limits_{0}^1\frac{\sqrt{1+e^{-x}}}{e^x}dx\]

OpenStudy (kainui):

Well, can we make any u-substitutions here? Looks like 1+e^-x might be a fruitful look if we see that 1/(e^x) is really e^-x.

terenzreignz (terenzreignz):

Here, I'll make it easier to see :P \[\Large \int\limits_{0}^1\frac{\sqrt{1+e^{-x}}}{e^x}dx = \int\limits_0^1e^{-x}\sqrt{1+e^{-x}}dx \]

OpenStudy (anonymous):

\[\Large \int\limits_{0}^1\frac{\sqrt{1+e^{-x}}}{e^x}dx =\Large \int\limits_{0}^1(\sqrt{1+e^{-x}})e^{-x}dx \]

OpenStudy (kainui):

Ok let's not completely spoon feed him now! lol

OpenStudy (anonymous):

Thanks for making it look viewable! So I'm pretty sure that u = e^-x so du = e^x dx

OpenStudy (anonymous):

one sec

terenzreignz (terenzreignz):

sorry.. wrong derivative...think chain rule :D

terenzreignz (terenzreignz):

And besides, that u-substitution won't help you... don't forget that you still have that constant term 1 in the radical that you have to worry about...

OpenStudy (anonymous):

Consider \(\sqrt{1+e^{-x}}=\sqrt{u(x)}\)

OpenStudy (anonymous):

-e^-x

terenzreignz (terenzreignz):

There you go... keep in mind that \(\large e^{-x}\) isn't the only function with \(\large -e^{-x}\) for its derivative....

OpenStudy (anonymous):

quick question, wouldn't i get a weird interval with this sub?

OpenStudy (anonymous):

like -1/e to 0 lol

terenzreignz (terenzreignz):

Where did you get that interval?

OpenStudy (anonymous):

u=1=-e^-1

OpenStudy (anonymous):

+1

terenzreignz (terenzreignz):

\[\Large \int\limits_a^b(f\circ g)(x)g'(x)dx \] Let u = g(x) du = g'(x)dx \[\Large \int\limits_{g(a)}^{g(b)}f(u)du \]

OpenStudy (anonymous):

Yes I understand that but the interval is still what I said right?

terenzreignz (terenzreignz):

Check it... if x = 0, what is \(\large 1+e^{-x}\)?

OpenStudy (kainui):

No, @iambatman you've chosen the wrong substitution. Pick: u=1+e^(-x) And hopefully you'll see why that works out well for you. It'll just take some practice to get the hang of "seeing" it.

OpenStudy (anonymous):

2

OpenStudy (anonymous):

ahh thanks kai lol

terenzreignz (terenzreignz):

and what about when x = 1? Work from there :D

OpenStudy (anonymous):

1+1/e?

terenzreignz (terenzreignz):

Yup, that's the one. I guess it is rather weird, but hey, that's the way things are :P

OpenStudy (anonymous):

I got that before and thought I was wrong since I've never seen anything like it before lol

OpenStudy (anonymous):

so I have a negative integral from 1+1e to 2 (square root (u)) du at the moment, so shall i go and get the anti deriv. of it now? :P

terenzreignz (terenzreignz):

negative integral from 1+1/e to 2? How?

OpenStudy (anonymous):

du = -e^-x dx, du/ -e^-x = dx and the intervals were changed to 2 and 1+1/e because u = 0 and u = 1

terenzreignz (terenzreignz):

Remember that THIS was your integral... \[\Large \int\limits_{0}^1(\sqrt{1+e^{-x}})e^{-x}dx\]

terenzreignz (terenzreignz):

Now, you let \(\large u = 1+e^{-x }\)

OpenStudy (anonymous):

yeah I did that du/dx = -e^-x

terenzreignz (terenzreignz):

Yes... -e^-x... something that you don't actually HAVE in your integrand...

OpenStudy (kainui):

Thank god for algebra, you can multiply both sides by -1. Disaster averted.

terenzreignz (terenzreignz):

To rectify that, we multiply it by (-1) two times... \[\Large \color{red}-\int\limits_{0}^1(\sqrt{1+e^{-x}})(\color{red} -e^{-x})dx\]

terenzreignz (terenzreignz):

Now, you may proceed normally.

OpenStudy (anonymous):

yeah thats what i have

terenzreignz (terenzreignz):

Then... continue it.

OpenStudy (kainui):

Show us what you're made of bat man. Help us help you by putting in the effort.

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

ok i'm getting some weird answer 4sqaureroot(2)/3 -2/3(1+1/e)^3/2

OpenStudy (anonymous):

Sweet, thanks for the help guys!